OpenStudy (anonymous):

solve 2x-y=1 if the domain of x is {-1,0,1}

OpenStudy (anonymous):

just substitute the equation with x = -1, 0 and 1 will do....

OpenStudy (anonymous):

2(-1)-y=1 multiply -2-y=1 add "2" 2 crosses out on the left on the rigt it becomes 3 -y=3 remove the negetive by dividing and y=-3 and do that so on and so forth withe the otuers plug in 0 and 1 for x

OpenStudy (anonymous):

I would probably start by solving for y: 2x-y=1 2x=1+y 2x-1 = y then plug in the values for x

OpenStudy (anonymous):

2(-1) - 1 = y -2 - 1 = y -3 = y 2(0) - 1 = y -1 = y 2(1) - 1 = y 2 - 1 = y 1 = y